Find the equivalent fraction of 3 / 5 having
denominator 20 

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

We require denominator 20
Let M be the numerator of the fractions
∴M / 20 = 3 / 5
5 × M = 20 × 3
M = (20 × 3) / 5
= 12
Therefore the required fraction is 12 / 20

Thanks